# On Developers’ Self-Declaration of AI-Generated Code: An Analysis of Practices

### v1

@@ 4.3. RQ1: How do developers self-declare AI-generated code in their projects?
2. Coded by Kristin Witte and GitHub Copilot (and ChatGPT) (comment)
@@ 5. Grayscale the entire page except the specified element solution by ChatGPT (comment)
7. Written by Brian Bird, 4/9/2024 with the assistance of GitHub Copilot (comment)
@@
Finding 3 Developers employ various methods to self-declare AI-generated code in their projects, including simply stating the use of AI in a single sentence, providing a brief explanation of the code, sharing contextual information (such as the prompts used to generate the code), or reporting the code quality either positively or negatively.
@@
Figure 13. An example of file-level self-declaration comment declaring the entire file as mostly AI-generated
@@
The responses to the industrial survey show similar results, with 33 participants (29.7%) indicating that they use file-level comments to self-declare their AI-generated code. A participant explained when they employ file-level comments: “I use file-level comments when I use intensively integrated AI-generated code, and I use snippet-level comments when integrate less AI-generated code” ($R_{10}$).
Snippet-level comments. Developers include these comments at specific locations of the source file (such as before a method/function or a class) to indicate that this particular portion of the code is AI-generated. For example, if a developer adds a comment above a function, it suggests that the comment pertains to that specific function. Figure 14 shows an example of snippet-level comment. In this example, the developer self-declared the AI-generated code before the fixIndentation() function. This suggests that the self-declaration comment applies only to the function. Developers self-declare various code sections in snippet-level comments, including methods/functions, classes, conditions, loops, or statements. Our findings indicate that developers primarily self-declare AI-generated code at the class and function level. We found 426 instances of snippet-level self-declaration comments, accounting for 69.4% of all collected comments.
In the industrial survey, 34 participants (30.6%) mentioned that they utilize snippet-level self-declaration comments to self-declare their AI-generated code, and reported the reasons such as showing exactly which part of the code is AI-generated as shared by $R_{19}$, “I use snippet-level comments to point out the AI-generated parts because it gives clear context right where it’s needed. This way, anyone looking at the code can easily see what was generated by AI”.
Figure 14. An example of snippet-level self-declaration comment declaring a function as AI-generated
Finding 4 Developers use snippet-level comments more often than file-level comments to self-declare AI-generated code. When the code file has a few AI-generated code snippets, developers prefer a snippet-level comment above each AI-generated code snippet. However, if the code is extensively AI-generated, they tend to write a single file-level comment at the beginning or end of the code file.
@@
We presented the distributions of AI-generated code in repositories and individual source code files in Figure 15.
@@
Finding 5 Most of the GitHub repositories in our collected dataset have a few code files with AI-generated code snippets. These code files with AI-generated code snippets typically contain a single self-declaration comment, with only a few code files having multiple self-declaration comments. Overall, the occurrences of self-declaration comments in GitHub repositories are low.
@@
Finding 6 Developers’ sentiments about the AI-generated code and tools used can be identified in some of the added self-declaration comments. The comments are mostly positive, acknowledging the usefulness of the AI code generation tools. However, developers also sometimes express negative sentiments, indicating their dissatisfaction with AI-generated code.
4.4. RQ2: Do developers self-declare AI-generated code in practice? Why or why not?
@@
Finding 7 Most participants revealed that they sometimes self-declare AI-generated code based on the proportion of AI-generated code included in the project and the extent of human modifications to the generated code.
@@
Tracking and monitoring code for later reviews and debugging Developers use self-declaration comments to track where AI-generated code is used in the project, allowing them to separate AI-generated and human-written code. This can be helpful for future reference, quality checks, debugging, and improvements. (28) 25.2%
@@
Evaluating and improving code quality Developers self-declare AI-generated code explaining the code quality, which shows the reliability of the code and helps in code improvement in the future, thus improving code readability. (15) 13.5%
@@
Tracking and monitoring code for later reviews and debugging. This is the most significant cited reason, as noted by 28 (25.2%) participants. Developers use self-declaration comments to track where the added AI-generated code is used in the project, which can be helpful for future reference, quality checks, debugging, or code review. By clearly self-declaring AI-generated code, developers can more easily trace and address any issues that arise in the future due to the AI-generated code. For example, a participant explained in $R_{9}$ that “by marking exactly where the AI-generated code starts and ends, it gives clear context to anyone reviewing or maintaining the code. They can easily differentiate between AI-generated and manually written parts without searching through the entire file”. Similarly, another participant mentioned that they “use AI-generated code to quickly create blocks of code that would be time-consuming to write. By declaring AI use, if issues arise later, it marks it as an area of code that is worth checking first”.
@@
Transparency, accountability, and ethical considerations. 23 (20.7%) participants noted reasons related to transparency, accountability, and ethics. These participants self-declare the origin of the code to maintain transparency in their teams as they believe that the ethical practice of using AI-generated code is to transparently self-declare it in the project. By including self-declaration comments, developers ensure clarity on which portions of code are AI-generated, promoting responsible use of AI and, thus, promoting accountability. For instance, a participant shared in $R_{108}$ self-declare AI-generated code to ensure transparency and maintain ethical standards, helping collaborators understand its origin and any unique aspects. This practice also supports clearer debugging and proper attribution”. Another participant mentioned in $R_{12}$self-declaring AI-generated code in projects promotes transparency and accountability. It helps users understand the source of the code, encourages responsible use, and allows for better collaboration between human developers and AI tools”.
Evaluating and improving code quality. 15 (13.5%) participants mentioned reasons related to documenting code quality for adding self-declaration comments. Added self-declaration comments are used to explain code quality, indicating the reliability of the code and aiding in future code improvement. For instance, it was noted that self-declaration comments “helps users understand the origin of the code, encourages responsible collaboration, and allows for better evaluation of the code’s quality and reliability” ($R_{27}$). Another participant also cited how self-declaration comments can provide an assessment of the code’s reliability: “it enables assessment of the code’s reliability based on testing and modification history” ($R_{39}$).
@@
Knowledge sharing within the team. 10 (9.0%) participants noted that they self-declare AI-generated code to encourage knowledge sharing within the team. When developers incorporate AI-generated code solutions into their projects, they help other project members understand the code by providing context for the AI-generated portion of the code. This promotes collaboration, helps other team members understand the code, and facilitates knowledge sharing among team members. A participant specifically emphasized sharing the prompts used to add context to the AI-generated code: “including the original prompt used to generate the AI code ensures context is provided for why the code was generated in that specific way” ($R_{71}$). Similarly, another participant reported that “the reason for self-declaring AI-generated code is to facilitate collaboration by providing context for other developers” ($R_{40}$).
@@
Finding 8 Most participants noted that they self-declare AI-generated code because of the need to track it for future code review and debugging due to concerns about code quality. Another reason that was mentioned was the promotion of the ethical use of AI-generated code.
@@
Customization and modification. One of the most frequently cited reasons, as noted by 13 (11.7%) participants, for not self-declaring AI-generated code is that developers mostly edit and thoroughly modify the generated code according to their specific use cases. It is believed that the AI-generated code may not fully address the business requirements or may contain inaccuracies. For these reasons, developers extensively verify, modify, and test the code. Since the code has undergone significant changes, developers feel that it has become their own contribution. For instance, a participant noted that he/she thoroughly modified the AI-generated code before committing it: “The AI-generated code that I use in my application is thoroughly understood, tested and, many times, radically changed before being committed, essentially making it original code” ($R_{93}$). Another participant pointed out that the AI-generated code does not fulfill requirements and emphasized review and modification in $R_{33}$-generated code does not fully meet business requirements and needs to be reviewed and modified”.
@@
Self-declaration is unnecessary. The other top-cited reason, as mentioned by 13 (11.7%) participants, for not self-declaring AI-generated code is that some developers see it as an unnecessary activity. It was noted that developers may utilize AI code generation tools to generate ideas; however, they emphasized that they only employ them to enhance their productivity. They see AI-generated code as similar to seeking help from online documentation or developer forums (such as Stack Overflow). Those AI code generation tools are seen as assisting tools. Hence, they believe that self-declaration of AI-generated code is an unnecessary activity. For example, a participant answered that: “I usually search for small code snippets in languages I don’t remember the syntax for, and I believe that adding comments specifying the source would be unnecessary for my team” ($R_{95}$). Another participant noted the following: “In the rare cases where I’m not providing my own original code, it’s going through multiple iterations of my prompting and manually editing to get what I need ….. I don’t self-declare using a spell-checker, so I don’t see the value in self-declaring AI-generated code”.
Finding 9 Participants noted that they do not self-declare AI-generated code because the code requires human review and modification before integration into their projects, and they feel that self-declaration is not required, as it is similar to using online documentation or tools like a spell checker.
@@ 5.1. Developers Practices in Declaring AI-Generated Code
Our study findings demonstrate that developers follow different ways to self-declare AI-generated code in their projects. Developers include varying levels of detail in the self-declaration comments, ranging from simple one-liners to detailed explanations and contextual information about the added code. Developers’ approaches to the scope of self-declaration comments also vary - sometimes self-declaring specific code snippets or entire code files as AI-generated. Developers utilize these ways not only to provide simple annotations but also to address ethical, collaborative, and quality-related concerns. Below, we discuss the key reasons why developers use various ways to self-declare AI-generated code.
5.1.1. Information Developers Mention in the Self-Declaration Comments
We classified the types of information provided by developers in the self-declaration comments, as shown below.
Simple self-declaration. Developers state in the self-declaration comments that the specific file or code snippet is AI-generated. Developers typically share the name of the tool used to generate the code, as well as the name of the code snippet generated (e.g., class or method name). This declaration approach is useful for code traceability as it separates AI-generated code from human-written code. Developers prefer writing simple self-declaration comments to maintain traceability and transparency when integrating AI-generated code while also focusing on efficiency when writing self-declaration comments to avoid complex workflows. Yu et al. also found that comments acknowledging the use of AI are mostly short and lack details (Yu et al., 2024). Simple self-declaration comments ensure that the development process remains transparent and adheres to ethical considerations. It also preserves the code provenance by documenting the origin and history of generated code.
Code explanation. Explanations of code are seen as useful, especially when the AI-generated code is relatively complex. Developers mostly write such comments to explain three aspects of the code: (1) its functionality, (2) how it was generated, and (3) how the code was added. By adding these comments, developers aim to enhance the comprehensibility and readability of the code, a goal also discussed in related studies (Pascarella et al., 2019; Nielebock et al., 2019; Scalabrino et al., 2016). Adding code explanations is particularly valuable for complex or critical code segments, where understanding the rationale behind the code is essential for debugging and future modifications. This way of self-declaration was also widely supported by the survey participants. However, the challenge with this self-declaration approach is to maintain a balance in the amount of details to be shared. This is because overly verbose explanations may hinder readability, while insufficient detail may fail to provide a meaningful explanation of the added AI-generated code.
Contextual information. Developers may also add contextual information in the self-declaration comments. For example, we noted several instances of comments that included the prompt used to generate the code, which can help inform other team members about the initial requirement for which the code was generated. This improves the team’s understanding of the code and its context. Developers also share other information about the code, including the date when the code was generated and the developer’s username for better tracking of AI-generated code. These results are also evident in our industrial survey, confirming that developers share contextual information when self-declaring AI-generated code. Using this practice enables developers to enhance knowledge sharing and collaboration within their teams. The study by Stalnaker et al. shows similar results, where developers mentioned that they record the prompts or note the whole interaction with AI tools to document AI-generated code (Stalnaker et al., 2024). Similarly, Yu et al. reported that developers share the prompts used to generate code (Yu et al., 2024).
Code quality indications. Self-declaration comments can contain code quality indications. This demonstrates developers’ awareness of the code quality risks associated with the use of AI-generated code. Developers often claim that they have integrated AI-generated code after human modification and testing before its use in their projects. On the contrary, they may also warn that the code offers a quick fix with less human intervention and inadequate testing. In both cases, developers mention details about the code quality to provide helpful information on the reliability of the integrated AI-generated code. This way of self-declaration is beneficial for addressing quality issues in AI-generated code, for example, AI-generated code may suffer from security issues (Fu et al., 2025) as well as bugs (Tambon et al., 2025).
5.1.2. Scope of Self-Declaration Comments
We categorized the scope of self-declaration comments into two categories: snippet-level and file-level self-declaration comments. We discuss both categories and their comparison below.
File-level comments. When developers add multiple AI-generated code snippets in one source file, or the source file is mainly written with assistance from AI code generation tools, they provide a single self-declaration comment at the start or end of the source file to acknowledge that the file contains AI-generated code. The drawback of file-level comments is that they are provided at a high level and can not help accurately track the AI-generated code (i.e., which part of the file is AI-generated).
Snippet-level comments. When developers use short AI-generated code snippets in their projects, they mention precisely which specific segments are AI-generated. The self-declaration comments only apply to the specific scope of the self-declared AI-generated code snippets. For example, if a developer writes a snippet-level comment above a function, then the scope of this comment applies only to that specific function. The main benefit of snippet-level self-declaration is that it accurately describes which code snippet is AI-generated, allowing for the separation of AI-generated from the human-written code within a code file. This helps in tracking the AI-generated code in later code reviews, bug fixing, and quality improvement processes.
@@ 5.1.2. Scope of Self-Declaration Comments
Snippet-level vs. file-level self-declaration comments. The preference between snippet-level and file-level self-declaration comments reflects developers’ trade-offs between granularity and efficiency. Developers prefer snippet-level self-declaration comments when they need to track and differentiate AI-generated code accurately. On the other hand, developers use file-level self-declaration comments when the code file contains multiple AI-generated code snippets or is predominantly AI-generated. File-level comments are efficient as they save time and effort, but sacrifice granularity, making it difficult to trace individual AI-generated code snippets within code files. Developers’ preference for snippet-level comments, as indicated by the industrial survey and mining study, demonstrates that they prioritize accurate traceability and accountability over efficiency when integrating AI-generated code in their projects.
5.1.3. Distribution of AI-Generated Code.
@@
5.1.4. Developer Satisfaction.
By analyzing developers’ sentiments expressed in their self-declaration comments, we found that most developers were satisfied and reported a positive experience in generating the desired AI-generated code. These results align with previous findings, which show that AI code generation tools enhance developers’ productivity (Peng et al., 2023) and are useful in coding tasks (Nikolaidis et al., 2024). These tools enable developers to focus on higher-level problem-solving tasks, as they reduce the time and effort required for development. However, this satisfaction is related to the ability to control and refine the outputs of code generation tools, as highlighted by the emphasis on iterative prompting and human intervention.

### v2

@@ 4.1. RQ1: How do developers self-declare AI-generated code in their projects?
2. Coded by [developer name] and GitHub Copilot (and ChatGPT) (comment)
@@ 5. Grayscale the entire page except the specified element solution by ChatGPT (comment)
7. Written by [developer name], 4/9/2024 with the assistance of GitHub Copilot (comment)
@@
Finding 1 Developers employ various methods to self-declare AI-generated code in their projects, including simply stating the use of AI in a single sentence, providing a brief explanation of the code, sharing contextual information (such as the prompts used to generate the code), or reporting the code quality either positively or negatively.
@@
Figure 11. Examples of file-level and snippet-level self-declaration comments
@@
The responses to the practitioners’ survey show similar results, with 33 participants (29.7%) indicating that they use file-level comments to self-declare their AI-generated code. A participant explained when they employ file-level comments: “I use file-level comments when I use intensively integrated AI-generated code, and I use snippet-level comments when integrate less AI-generated code” ($R_{10}$).
Snippet-level comments. Developers include these comments at specific locations of the source file (such as before a method/function or a class) to indicate that this particular portion of the code is AI-generated. For example, if a developer adds a comment above a function, it suggests that the comment pertains to that specific function. Figure 11 shows an example of snippet-level comment. In this example, the developer self-declared the AI-generated code before the fixIndentation() function. This suggests that this self-declaration comment applies only to the function. Developers self-declare various code sections in snippet-level comments, including methods/functions, classes, conditions, loops, or statements. Our findings indicate that developers primarily self-declare AI-generated code at the class and function level. We found 426 instances of snippet-level self-declaration comments, accounting for 69.4% of all collected comments.
Figure 12. LoC of files containing file-level and snippet-level self-declaration comments
In the practitioners’ survey, 34 participants (30.6%) mentioned that they utilize snippet-level self-declaration comments to self-declare their AI-generated code, and reported the reasons such as showing exactly which part of the code is AI-generated as shared by $R_{19}$, “I use snippet-level comments to point out the AI-generated parts because it gives clear context right where it’s needed. This way, anyone looking at the code can easily see what was generated by AI”.
We also calculated the lines of code (LoC) of all 613 code files in our dataset. Then, we compared the LoC of files containing snippet-level comments with files containing file-level comments to understand their relationship with file size. Figure 12 demonstrates that snippet-level comments are found in larger code files compared to file-level comments found in shorter code files. The statistical analysis of LoC data reveals a substantial difference in file sizes: files with snippet-level comments have a mean of 436 LoC (median 215), while files with file-level comments average 190 LoC (median 86). Although there is high variability in file sizes for snippet-level comments, 54% of the files are over 200 LoC. On the contrary, 75% of the files containing file-level comments are below 200 lines of code. This significant difference in the size of files with snippet-level and file-level comments suggests that developers tend to use snippet-level comments for larger, more complex files and file-level comments for shorter code files.
Finding 2 Developers use snippet-level comments more often than file-level comments to self-declare AI-generated code. When the code file has a few AI-generated code snippets, developers prefer a snippet-level comment above each AI-generated code snippet. However, if the code is extensively AI-generated, they tend to write a single file-level comment at the beginning or end of the code file.
@@
We presented the distributions of AI-generated code in repositories and individual source code files in Figure 13. In this section, we refer specifically to self-declared AI-generated code, acknowledging that the actual proportion of AI-generated code in repositories may be higher.
@@
Figure 14. Monthly trend of self-declared AI-generated code spanning Jan 2021 to December 2024
Trend of self-declaring AI-generated code among developers. We visualized the dates when self-declaration comments were added to the files, covering the period from January 2021 to December 2024. We observed an increasing trend in self-declaring AI-generated code among developers, as shown in Figure 14. The trend exhibits continuous growth, punctuated by sudden spikes. A noticeable increase is observed starting in November 2022, coinciding with the introduction of the GPT-3.5 model. Similarly, we observed another significant rise in March 2023, when the GPT-4 model was launched alongside its integration with Copilot X. Additionally, there was a further spike in November 2023, following the release of GPT-4 turbo. Overall, the trend diagram illustrates that the practice of self-declaring AI-generated code is on the rise as more advanced models are introduced. However, we see a sudden drop in the trend after April 2024, which may be attributed to the timing of our data collection from GitHub. Consequently, newer self-declared AI-generated code snippets may not have been fully indexed on GitHub yet or may not have gained sufficient visibility to be captured in our analysis.
Finding 3 Most of the GitHub repositories in our collected dataset have a few code files with AI-generated code snippets. These code files with AI-generated code snippets typically contain a single self-declaration comment, with only a few code files having multiple self-declaration comments. Overall, the occurrences of self-declaration comments in GitHub repositories are low. The trend indicates a steady increase in self-declaring AI-generated code among developers.
@@
We also found reasons for developers’ satisfaction that they shared in their positive self-declaration comments. The main reasons we identified are that AI code generation tools save developers time and provide working solutions, help them solve complex logical challenges, assist them in improving the code, and identify critical errors in code. For example, a developer appreciated ChatGPT for helping them generate working code in less time, “Btw, everything in class LatencyMonitor (including the class and method docstring comments) was written by ChatGPT. It took about 10-15 minutes of interactive code review between me and ChatGPT to get to this point. Amazing!” (comment). Similarly, a developer acknowledged the help of ChatGPT in identifying the problem in code and improving it, “makes sure the user only sees relevant options (only for languages so far), note: ChatGPT helped make this way more efficient - was querying the db too much” (comment).
@@
We also identified reasons for developers’ dissatisfaction shared in their negative self-declaration comments. The main reason for their dissatisfaction is the poor code quality of the AI-generated code. The code quality issues are related to readability and maintainability, code optimization and efficiency, and error handling. For instance, a developer shared that they do not trust the code generated by Copilot because it produced inefficient code, “I wrote this with Github Copilot, don’t trust it … Probably uses tons of API calls. TODO: Can probably be optimized by first getting all the issuesPRscomments, putting them in some cache, and then iterating over them” (comment). Similarly, another developer indicated code readability and maintainability issues in the Copilot-generated code, “Please excuse the comments and lengthy code, this was written using Copilot and I have not yet cleaned it up” (comment).
Finding 4 Developers’ sentiments about the AI-generated code and tools used can be identified in some of the added self-declaration comments. The comments are mostly positive, acknowledging the usefulness of the AI code generation tools. However, developers also sometimes express negative sentiments, indicating their dissatisfaction with AI-generated code.
4.2. RQ2: What proportion of developers self-declare AI-generated code in practice? Why or why not?
@@
Finding 5 Most participants revealed that they sometimes self-declare AI-generated code based on the proportion of AI-generated code included in the project and the extent of human modifications to the generated code.
@@
Tracking and monitoring code for later reviews and debugging Developers use self-declaration comments to track where AI-generated code is used in the project, allowing them to separate AI-generated and human-written code. This can be helpful for future reference, quality checks, debugging, and improvements. (43) 38.7%
@@
Tracking and monitoring code for later reviews and debugging. This is the most significant cited reason, as noted by 43 (38.7%) participants. Developers use self-declaration comments to track where the added AI-generated code is used in the project, which can be helpful for future reference, quality checks, debugging, or code review. By clearly self-declaring AI-generated code, developers can more easily trace and address any issues that arise in the future due to the AI-generated code. For example, a participant explained in $R_{9}$ that self-declaration is useful for differentiating AI-generated code from human-authored code, “by marking exactly where the AI-generated code starts and ends, it gives clear context to anyone reviewing or maintaining the code. They can easily differentiate between AI-generated and manually written parts without searching through the entire file”. Similarly, another participant mentioned in $R_{109}$ that they “use AI-generated code to quickly create blocks of code that would be time-consuming to write. By declaring AI use, if issues arise later, it marks it as an area of code that is worth checking first”. Another participant shared that they self-declare the code when it extensively integrates AI-generated code in $R_{10}$, “I always self-declare the code when it’s extensively AI-generated. It’s because we need to know which part of the code is AI-generated for testing and code review purposes. But, if the code just took an idea from the AI-gen tool and then code is written by human modification as per the requirements [sic] I optionally self-declare”. Some participants also shared that self-declaring AI-generated code is also helpful in recording the code quality as in $R_{04}$, “It helps to keep a record of code quality and clarifies the proportion of AI-generated code versus human-written code. It also demonstrates that the code has been fully tested”.
@@
Transparency, accountability, and ethical considerations. 23 (20.7%) participants noted reasons related to transparency, accountability, and ethics. These participants self-declare the origin of the code to maintain transparency in their teams as they believe that the ethical practice of using AI-generated code is to transparently self-declare it in the project. By including self-declaration comments, developers ensure clarity on which portions of code are AI-generated, promoting responsible use of AI and, thus, promoting accountability. For instance, a participant shared in $R_{108}$I self-declare AI-generated code to ensure transparency and maintain ethical standards, helping collaborators understand its origin and any unique aspects. This practice also supports clearer debugging and proper attribution”. Another participant mentioned in $R_{12}$self-declaring AI-generated code in projects promotes transparency and accountability. It helps users understand the source of the code, encourages responsible use, and allows for better collaboration between human developers and AI tools”. A participant particularly emphasized accountability when using AI-generated code in $R_{66}$, “Accountability: Declaring the tool ensures that the use of AI in coding is fully documented and traceable”.
@@
Knowledge sharing within the team. 10 (9.0%) participants noted that they self-declare AI-generated code to encourage knowledge sharing within the team. When developers incorporate AI-generated code solutions into their projects, they help other project members understand the code by providing context for the AI-generated portion of the code. This promotes collaboration, helps other team members understand the code, and facilitates knowledge sharing among team members. A participant specifically emphasized sharing the prompts used to add context to the AI-generated code: “including the original prompt used to generate the AI code ensures context is provided for why the code was generated in that specific way” ($R_{71}$). Similarly, another participant reported that “the reason for self-declaring AI-generated code is to facilitate collaboration by providing context for other developers” ($R_{40}$). A participant expressed that they self-declare AI-generated code for the understanding of other developers in the team in ($R_{72}$), “Other developers can understand the logic or task behind the AI-generated code, making it easier to validate or modify it”.
@@
Finding 6 Most participants noted that they self-declare AI-generated code because of the need to track it for future code review and debugging due to concerns about code quality. Another reason that was mentioned was the promotion of the ethical use of AI-generated code.
@@
Customization and modification. One of the most frequently cited reasons, as noted by 13 (11.7%) participants, for not self-declaring AI-generated code is that developers mostly edit and thoroughly modify the generated code according to their specific use cases. It is believed that the AI-generated code may not fully address the business requirements or may contain inaccuracies. For these reasons, developers extensively verify, modify, and test the code. Since the code has undergone significant changes, developers feel that it has become their own contribution. For instance, a participant noted that they thoroughly modify the AI-generated code before committing it: “The AI-generated code that I use in my application is thoroughly understood, tested and, many times, radically changed before being committed, essentially making it original code” ($R_{93}$). Another participant shared a similar idea but also expressed that they can trust the AI-generated code as their own code after their manual review and modification, “I review code thoroughly and modify it if it’s required, so it’s no longer AI-generated code. I can trust it as much as I trust myself” ($R_{91}$). In addition, a participant pointed out that the AI-generated code does not fulfill requirements in $R_{33}$AI-generated code does not fully meet business requirements and needs to be reviewed and modified”. Some participants emphasized that they use AI-generated code to take an initial idea to save time, but they do not use this code blindly, as explained in the next two responses. In $R_{110}$, the developer mentioned that “It’s irrelevant where the code came from. I don’t blindly paste code and move on, I read it, test it, and comprehend it, like any other code I would write”. Similarly, a developer shared in $R_{03}$ that “I only took an initial idea then modify [sic] it to meet my requirements for the problem I work on and then use it. This helps me to not search for docs and saves my time. I also add comments explaining the code if necessary”.
@@
Self-declaration is unnecessary. The other top-cited reason, as mentioned by 13 (11.7%) participants, for not self-declaring AI-generated code is that some developers see it as an unnecessary activity. It was noted that developers may utilize AI code generation tools to generate ideas; however, they emphasized that they only employ them to enhance their productivity. They see AI-generated code as similar to seeking help from online documentation or developer forums (such as Stack Overflow). Those AI code generation tools are seen as assisting tools. Hence, they believe that self-declaration of AI-generated code is an unnecessary activity. For example, a participant answered that: “I usually search for small code snippets in languages I don’t remember the syntax for, and I believe that adding comments specifying the source would be unnecessary for my team” ($R_{95}$). Another participant noted the following: “In the rare cases where I’m not providing my own original code, it’s going through multiple iterations of my prompting and manually editing to get what I need ….. I don’t self-declare using a spell-checker, so I don’t see the value in self-declaring AI-generated code”. A participant expressed that they use AI code generation tools as an assisting tool but not as an autonomous contributor in $R_{92}$, “I use Github Copilot, which writes code alongside me. When I am doing any functionality, it keeps giving me suggestions and one of them suits me so I just select it and then after making desirable changes, it’s good to go. Since it’s combination of both me and Copilot writing code, I don’t feel the need of specifying something”. Some participants mentioned that they see AI-generated code similar to finding solutions from code documentation or developer forums, as in $R_{87}$No need [for self-declaring AI-generated code] because it would be like declaring some code that comes from documentation or StackOverflow or a colleagues’ [sic] advice”.
Finding 7 Participants noted that they do not self-declare AI-generated code because the code requires human review and modification before integration into their projects, and they feel that self-declaration is not required, as it is similar to using online documentation or tools like a spell checker.
@@ 5.1. Developers Practices in Declaring AI-Generated Code
In this section, we discuss the key reasons why developers use various ways to self-declare AI-generated code.
5.1.1. Developers Mostly Use Simple Self-Declaration Comments.
In our mining study, the highest number of self-declaration comments were simple one-line comments, indicating that the code was generated by an AI code generation tool. This finding aligns with our practitioners’ survey results, where the top reasons for self-declaring AI-generated code were tracking and monitoring AI-generated code and fulfilling ethical or transparency requirements. These goals can be effectively met using simple self-declaration comments, which allow developers to mark AI-generated code without adding additional information that may make self-declaration inefficient. Therefore, these results from the mining study and the practitioners’ survey complement each other: simple self-declaration comments enable both accurate traceability and ethical acknowledgment of AI use in a practical and straightforward manner. These findings align with existing studies, such as Yu et al., who also found that comments acknowledging the use of AI are mostly short and lack details (Yu et al., 2024). Additionally, simple self-declaration comments help developers maintain transparency, preserve the provenance of generated code, and integrate AI-generated code efficiently without adding unnecessary complexity to their workflows.
5.1.2. Developers Attach Additional Information in Self-Declaration Comments.
Our mining study findings show that developers also enrich self-declaration comments with contextual information, such as prompts, code explanations, and code quality indicators. Developers utilize these ways not only to provide simple annotations but also to address ethical, collaborative, and quality-related concerns on AI-generated code. These ways of self-declaring AI-generated code have various benefits for developers. (1) Self-declaration comments can contain code quality indications sharing details about the quality and the reliability of the integrated AI-generated code, which is also highlighted in existing literature (Fu et al., 2025; Tambon et al., 2025). (2) Self-declaration comments can also contain contextual information or code explanations explaining the context and rationale of integrated AI-generated code. This practice is also aligned with existing studies, such as the survey by Stalnaker et al., which shows that developers record the prompts or note the whole interaction with AI tools to document AI-generated code (Stalnaker et al., 2024). Our survey participants also supported these ways of writing self-declaration comments (see Figure 10). However, the challenge with attaching additional information is to maintain a balance in the amount of details to be shared. This is because overly detailed comments may hinder readability, while insufficient detail may fail to provide meaningful information about the added AI-generated code.
5.1.3. Snippet-level vs. File-level self-Declaration Comments.
@@
The preference between snippet-level and file-level self-declaration comments reflects developers’ trade-offs between granularity and efficiency. Developers prefer snippet-level self-declaration comments when they need to track and differentiate AI-generated code accurately. On the other hand, developers use file-level self-declaration comments when the code file contains code that is predominantly AI-generated. File-level comments are efficient as they save time and effort, but sacrifice granularity, making it difficult to trace individual AI-generated code snippets within code files. Developers’ preference for snippet-level comments, as indicated by the practitioners’ survey and mining study, demonstrates that they prioritize accurate traceability and accountability over efficiency when integrating AI-generated code in their projects.
5.1.4. Most Repositories Contain Only a Single File With Self-Declared AI-Generated Code.
While 69% of repositories in the mining study contained only one file with self-declared AI-generated code, our survey found that 63% of developers “sometimes” self-declare AI-generated code. At first glance, this might appear inconsistent, but a closer look clarifies the difference. The survey results indicate that these developers self-declare AI-generated code selectively, in cases when the code has minimal human intervention or has not been properly tested. It can be explained using a participant’s response in $R_{51}$, “I rarely use it in some part of my project so I only declare when it is used more than 30% in my project”. Only 13.5% of developers reported “always” self-declaring AI-generated code. This selective behavior explains why repositories often contain only a single file with such comments: the majority of developers self-declare AI-generated code only in selective cases when they consider necessary.
Furthermore, the survey results suggest that developers are generally willing to self-declare AI-generated code, with 63% indicating that they self-declare AI-generated code at least “sometimes”, and only 23% reporting that they never self-declare AI-generated code. This willingness implies that with better tool support, such as automated mechanisms for detecting AI-generated code and helping developers insert self-declaration comments, the adoption of the practices of self-declaring AI-generated code could increase significantly. This is likely because these tools would reduce the perceived effort and liability associated with self-declaring AI-generated code.
5.1.5. Distribution of Self-Declared AI-Generated Code in GitHub Repositories.